# Branching Brownian motion versus Random Energy Model in the supercritical phase: overlap distribution and temperature susceptibility

### v2

@@ 3.1 Results in the general decorated case
Part (ii). We check again that the assumptions of Theorem 3.1 are satisfied. Since $\psi$ is linear in a neighborhood of 1, we have $\psi(1\pm h)=\psi(1)\pm h\psi^{\prime}(1)$ for $h$ small enough and Equation (3.5) implies $\mathbb{E}[T_{\beta}\log S_{\beta}^{1/\beta}]=\psi^{\prime}(1)\log\frac{1}{\beta-1}+O(1)$. Combining this with Equation (3.4), we get Assumption (i’) with $\alpha=\psi^{\prime}(1)-\psi(1)=0$. Assumption (ii) follows from Equation (3.6) and the fact that $\psi(1+\varepsilon)-(1+\varepsilon)\psi(1)=\varepsilon(\psi^{\prime}(1)-\psi(1))=\varepsilon$, and Assumption (iii) is obtained as before. Hence, Theorem 3.1 implies $F_{d}(\beta,\beta^{\prime})=O(\beta-1)$. But under these assumptions Theorem 3.3 can also be applied and yield the desired lower bound.
@@
by our choice of $\gamma$. Hence, we can apply Theorem 3.4 and get $F_{d}(\beta,\beta^{\prime})=O((\beta-1)^{1+\eta})$, which proves the result. ∎
@@
We introduce a family of decoration processes $(\mathcal{D}^{a,b})_{a,b>0}$ and study the behavior of $F_{d}(\beta,\beta^{\prime})$ for these cases. For $a,b>0$, let $X_{a}\geq 1$ be a random variable with law defined by
@@
and let $\mathcal{D}^{a,b}$ be a point process such that, conditionally on $X_{a}$, $\mathcal{D}^{a,b}$ is the sum of a Dirac mass at 0 and a Poisson point process with intensity $\mathopen{}\mathclose{{}\left\lvert x}\right\rvert^{b-1}\mathrm{e}^{-x}\mathbbm{1}_{x\in[-X_{a},0]}\mathop{}\mathopen{}\mathrm{d}x$. This decoration is a toy model of the decoration of the BBM when $a=1$ and $b=2$, see Remarks 2.9 and 2.11. We are going to check that, for any $\beta^{\prime}>1$, as $\beta\downarrow 1$,
@@
Note that one does not treat the case $a=b$, which do not fits in the theorems of this section. In the last case, by taking for example $b=2a$ large enough, one can have $F_{d}(\beta,\beta^{\prime})$ vanishing faster than any power of $(\beta-1)$.
@@ 3.2 Some tools for the proofs
The starting point for the proof of the results stated in the previous section is the following expression for $F_{d}(\beta,\beta^{\prime})$.
Lemma 3.7.
@@
$R=R(\beta,\beta^{\prime})\coloneqq\biggl{(}\frac{Z_{d}(\beta)}{S_{\beta}}\biggr{)}^{1/\beta}\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}.$
@@
$F_{d}(\beta,\beta^{\prime})=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\int_{0}^{\infty}\frac{\mathop{}\mathopen{}\mathrm{d}x}{(1+(Rx)^{\beta})(1+x^{\beta^{\prime}})}}\right].$
@@
$F(\beta,\beta^{\prime})$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\sum_{k\geq 1}\eta_{k}^{-(\beta+\beta^{\prime})}S_{\beta,k}S_{\beta^{\prime},k}\frac{1}{\Bigl{(}\eta_{k}^{-\beta}S_{\beta,k}+\sum_{j\neq k}\eta_{j}^{-\beta}S_{\beta,j}\Bigr{)}\Bigl{(}\eta_{k}^{-\beta^{\prime}}S_{\beta^{\prime},k}+\sum_{j\neq k}\eta_{j}^{-\beta^{\prime}}S_{\beta^{\prime},j}\Bigr{)}}}\right]$
$=\int_{(\mathbb{R}_{+})^{3}}\mathbb{E}\mathopen{}\mathclose{{}\left[x^{-(\beta+\beta^{\prime})}ss^{\prime}\frac{1}{(x^{-\beta}s+Z_{d}(\beta))(x^{-\beta^{\prime}}s^{\prime}+Z_{d}(\beta^{\prime}))}}\right]\mathop{}\mathopen{}\mathrm{d}x\mathop{}\mathopen{}\mathrm{d}\nu(s,s^{\prime}),$
where $\nu$ denotes the law of $(S_{\beta},S_{\beta^{\prime}})$ and one applied Palm formula (see Proposition 3.8 stated below) to the Poisson point process $\sum_{i}\delta_{(\eta_{i},S_{\beta,i},S_{\beta^{\prime},i})}$ on $(\mathbb{R}_{+})^{3}$, which has intensity $\mathop{}\mathopen{}\mathrm{d}x\otimes\mathop{}\mathopen{}\mathrm{d}\nu(s,s^{\prime})$. Using Fubini’s theorem together with the fact that $(S_{\beta},S_{\beta^{\prime}})$ is independent of $(Z_{d}(\beta),Z_{d}(\beta^{\prime}))$ yields
$F_{d}(\beta,\beta^{\prime})=\mathbb{E}\mathopen{}\mathclose{{}\left[\int_{0}^{\infty}\frac{\mathop{}\mathopen{}\mathrm{d}x}{(1+x^{\beta}Z_{d}(\beta)/S_{\beta})(1+x^{\beta^{\prime}}Z_{d}(\beta^{\prime})/S_{\beta^{\prime}})}}\right].$
@@ 3.3 Proof of the general theorems
$F_{d}(\beta,\beta^{\prime})$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\frac{\log R}{R}}\right]+O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\mathopen{}\mathclose{{}\left(\frac{1}{R}+\frac{(\beta-1)}{R^{1-\delta}}+\frac{1}{R^{1+\delta}}}\right)}\right]}\right)$
@@
Term $E_{1}$. Using the definition of $R$ and then Corollary 2.7 together with the independence of $S_{\beta}$ and $Z_{d}(\beta)$, we get
$E_{1}=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{d}(\beta)}}\right)^{1/\beta}}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-1/\beta}}\right]\sim\beta-1,\qquad\beta\downarrow 1,$ (3.10)
@@
Term $E_{3}$. We use the definition of $R$ and the independence of $(S_{\beta},S_{\beta^{\prime}})$ and $(Z_{d}(\beta),Z_{d}(\beta^{\prime}))$, and then $S_{\beta^{\prime}}\geq 1$ and Cauchy–Schwarz inequality to get
$E_{3}$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{S_{\beta}^{(1+\delta)/\beta}}{S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{d}(\beta^{\prime})^{\delta/\beta^{\prime}}}{Z_{d}(\beta)^{(1+\delta)/\beta}}}\right]\leq\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1+\delta)/\beta}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[Z_{d}(\beta^{\prime})^{2\delta/\beta^{\prime}}}\right]^{1/2}\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1}{Z_{d}(\beta)^{2(1+\delta)/\beta}}}\right]^{1/2}.$
Using Corollary 2.7, Lemma 2.2 and $\mathbb{E}[S_{\beta^{\prime}}^{1/\beta^{\prime}}]<\infty$, we get $\mathbb{E}[Z_{d}(\beta^{\prime})^{2\delta/\beta^{\prime}}]=O(1)$ as soon as $\delta<1/2$. Therefore, by Corollary 2.7 and Equation (2.3),
@@
$E_{2}$ $=(\beta-1)\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\delta)/\beta}S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{d}(\beta^{\prime})^{-\delta/\beta^{\prime}}}{Z_{d}(\beta)^{(1-\delta)/\beta}}}\right]=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\delta)/\beta}S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}\right]\cdot\frac{(\beta-1)^{2-\delta}}{\mathbb{E}[S_{\beta}^{1/\beta}]^{1-\delta}}}\right).$
@@
$E_{0}$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{d}(\beta)}}\right)^{1/\beta}\mathopen{}\mathclose{{}\left(\log\frac{Z_{d}(\beta)^{1/\beta}}{S_{\beta}^{1/\beta}}+\frac{1}{\beta^{\prime}}\log S_{\beta^{\prime}}-\frac{1}{\beta^{\prime}}\log Z_{d}(\beta^{\prime})}\right)}\right]\eqqcolon E_{00}+E_{01}-E_{02},$
@@
$E_{02}$ $=\frac{1}{\beta^{\prime}}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{\log Z_{d}(\beta^{\prime})}{Z_{d}(\beta)^{1/\beta}}}\right]=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\log^{2}\mathopen{}\mathclose{{}\left(Z(\beta^{\prime})\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]}\right)}\right]^{1/2}\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-2/\beta}}\right]^{1/2}}\right),$
@@
$F_{d}(\beta,\beta^{\prime})\geq\frac{1}{4}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\mathopen{}\mathclose{{}\left(1\wedge\frac{1}{R}}\right)}\right]=\frac{1}{4}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\wedge\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{d}(\beta)}}\right)^{1/\beta}}\right].$ (3.12)
@@
$F_{d}(\beta,\beta^{\prime})\geq\frac{1}{4M}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{d}(\beta)}}\right)^{1/\beta}}\right]-\frac{1}{M^{\delta}}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{Z_{d}(\beta^{\prime})}{S_{\beta^{\prime}}}}\right)^{\delta/\beta^{\prime}}\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{d}(\beta)}}\right)^{(1+\delta)/\beta}}\right]}\right).$ (3.13)
@@
$F_{d}(\beta,\beta^{\prime})=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{d}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\frac{1}{R^{1-\theta}}}\right]}\right)=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\theta)/\beta}S_{\beta^{\prime}}^{\theta/\beta^{\prime}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{d}(\beta^{\prime})^{-\theta/\beta^{\prime}}}{Z_{d}(\beta)^{(1-\theta)/\beta}}}\right]}\right),$ (3.14)
using the definition of $R$ and the independence of $(S_{\beta},S_{\beta^{\prime}})$ and $(Z_{d}(\beta),Z_{d}(\beta^{\prime}))$. By Cauchy–Schwarz inequality and Corollary 2.7,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{d}(\beta^{\prime})^{-\theta/\beta^{\prime}}}{Z_{d}(\beta)^{(1-\theta)/\beta}}}\right]\leq\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta^{\prime})^{-2\theta/\beta^{\prime}}}\right]^{1/2}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]^{\theta}}\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-2(1-\theta)/\beta}}\right]^{1/2}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{1-\theta}}=O\mathopen{}\mathclose{{}\left(\frac{(\beta-1)^{1-\theta}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]^{\theta}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{1-\theta}}}\right),$

### v3

@@ 3.1 Results in the general decorated case
Part (ii). We check again that the assumptions of Theorem 3.1 are satisfied. Since $\psi$ is linear in a neighborhood of 1, we have $\psi(1\pm h)=\psi(1)\pm h\psi^{\prime}(1)$ for $h$ small enough and Equation (3.5) implies $\mathbb{E}[T_{\beta}\log S_{\beta}^{1/\beta}]=\psi^{\prime}(1)\log\frac{1}{\beta-1}+O(1)$. Combining this with Equation (3.4), we get Assumption (i’) with $\alpha=\psi^{\prime}(1)-\psi(1)=0$. Assumption (ii) follows from Equation (3.6) and the fact that $\psi(1+\varepsilon)-(1+\varepsilon)\psi(1)=\varepsilon(\psi^{\prime}(1)-\psi(1))=\varepsilon$, and Assumption (iii) is obtained as before. Hence, Theorem 3.1 implies $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=O(\beta-1)$. But under these assumptions Theorem 3.3 can also be applied and yield the desired lower bound.
@@
by our choice of $\gamma$. Hence, we can apply Theorem 3.4 and get $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=O((\beta-1)^{1+\eta})$, which proves the result. ∎
@@
We introduce a family of decoration processes $(\mathcal{D}^{a,b})_{a,b>0}$ and study the behavior of $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ for these cases. For $a,b>0$, let $X_{a}\geq 1$ be a random variable with law defined by
@@
and let $\mathcal{D}^{a,b}$ be a point process such that, conditionally on $X_{a}$, $\mathcal{D}^{a,b}$ is the sum of a Dirac mass at 0 and a Poisson point process with intensity $\mathopen{}\mathclose{{}\left\lvert x}\right\rvert^{b-1}\mathrm{e}^{-x}\mathbbm{1}_{x\in[-X_{a},0]}\mathop{}\mathopen{}\mathrm{d}x$. This decoration is a toy model of the decoration of the BBM when $a=1$ and $b=2$, see Remarks 2.10 and 2.12. We are going to check that, for any $\beta^{\prime}>1$, as $\beta\downarrow 1$,
@@
Note that one does not treat the case $a=b$, which do not fits in the theorems of this section. In the last case, by taking for example $b=2a$ large enough, one can have $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ vanishing faster than any power of $(\beta-1)$.
@@ 3.2 Some tools for the proofs
The starting point for the proof of the results stated in the previous section is the following expression for $\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$.
Lemma 3.7 (General decorated case).
@@
$R=R(\beta,\beta^{\prime})\coloneqq\biggl{(}\frac{Z_{\mathrm{d}}(\beta)}{S_{\beta}}\biggr{)}^{1/\beta}\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}.$
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\int_{0}^{\infty}\frac{\mathop{}\mathopen{}\mathrm{d}x}{(1+(Rx)^{\beta})(1+x^{\beta^{\prime}})}}\right].$
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\sum_{k\geq 1}\eta_{k}^{-(\beta+\beta^{\prime})}S_{\beta,k}S_{\beta^{\prime},k}\frac{1}{\Bigl{(}\eta_{k}^{-\beta}S_{\beta,k}+\sum_{j\neq k}\eta_{j}^{-\beta}S_{\beta,j}\Bigr{)}\Bigl{(}\eta_{k}^{-\beta^{\prime}}S_{\beta^{\prime},k}+\sum_{j\neq k}\eta_{j}^{-\beta^{\prime}}S_{\beta^{\prime},j}\Bigr{)}}}\right]$
$=\int_{(\mathbb{R}_{+})^{3}}\mathbb{E}\mathopen{}\mathclose{{}\left[x^{-(\beta+\beta^{\prime})}ss^{\prime}\frac{1}{(x^{-\beta}s+Z_{\mathrm{d}}(\beta))(x^{-\beta^{\prime}}s^{\prime}+Z_{\mathrm{d}}(\beta^{\prime}))}}\right]\mathop{}\mathopen{}\mathrm{d}x\mathop{}\mathopen{}\mathrm{d}\nu(s,s^{\prime}),$
where $\nu$ denotes the law of $(S_{\beta},S_{\beta^{\prime}})$ and one applied Palm formula (see Proposition 3.8 stated below) to the Poisson point process $\sum_{i}\delta_{(\eta_{i},S_{\beta,i},S_{\beta^{\prime},i})}$ on $(\mathbb{R}_{+})^{3}$, which has intensity $\mathop{}\mathopen{}\mathrm{d}x\otimes\mathop{}\mathopen{}\mathrm{d}\nu(s,s^{\prime})$. Using Fubini’s theorem together with the fact that $(S_{\beta},S_{\beta^{\prime}})$ is independent of $(Z_{\mathrm{d}}(\beta),Z_{\mathrm{d}}(\beta^{\prime}))$ yields
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[\int_{0}^{\infty}\frac{\mathop{}\mathopen{}\mathrm{d}x}{(1+x^{\beta}Z_{\mathrm{d}}(\beta)/S_{\beta})(1+x^{\beta^{\prime}}Z_{\mathrm{d}}(\beta^{\prime})/S_{\beta^{\prime}})}}\right].$
@@ 3.3 Proof of the general theorems
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\frac{\log R}{R}}\right]+O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\mathopen{}\mathclose{{}\left(\frac{1}{R}+\frac{(\beta-1)}{R^{1-\delta}}+\frac{1}{R^{1+\delta}}}\right)}\right]}\right)$
@@
Term $E_{1}$. Using the definition of $R$ and then Corollary 2.8 together with the independence of $S_{\beta}$ and $Z_{\mathrm{d}}(\beta)$, we get
$E_{1}=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{\mathrm{d}}(\beta)}}\right)^{1/\beta}}\right]=\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-1/\beta}}\right]\sim\beta-1,\qquad\beta\downarrow 1,$ (3.10)
@@
Term $E_{3}$. We use the definition of $R$ and the independence of $(S_{\beta},S_{\beta^{\prime}})$ and $(Z_{\mathrm{d}}(\beta),Z_{\mathrm{d}}(\beta^{\prime}))$, and then $S_{\beta^{\prime}}\geq 1$ and Cauchy–Schwarz inequality to get
$E_{3}$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{S_{\beta}^{(1+\delta)/\beta}}{S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{\mathrm{d}}(\beta^{\prime})^{\delta/\beta^{\prime}}}{Z_{\mathrm{d}}(\beta)^{(1+\delta)/\beta}}}\right]\leq\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1+\delta)/\beta}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[Z_{\mathrm{d}}(\beta^{\prime})^{2\delta/\beta^{\prime}}}\right]^{1/2}\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1}{Z_{\mathrm{d}}(\beta)^{2(1+\delta)/\beta}}}\right]^{1/2}.$
Using Corollary 2.8, Lemma 2.3 and $\mathbb{E}[S_{\beta^{\prime}}^{1/\beta^{\prime}}]<\infty$, we get $\mathbb{E}[Z_{\mathrm{d}}(\beta^{\prime})^{2\delta/\beta^{\prime}}]=O(1)$ as soon as $\delta<1/2$. Therefore, by Corollary 2.8 and Equation (2.3),
@@
$E_{2}$ $=(\beta-1)\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\delta)/\beta}S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{\mathrm{d}}(\beta^{\prime})^{-\delta/\beta^{\prime}}}{Z_{\mathrm{d}}(\beta)^{(1-\delta)/\beta}}}\right]=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\delta)/\beta}S_{\beta^{\prime}}^{\delta/\beta^{\prime}}}\right]\cdot\frac{(\beta-1)^{2-\delta}}{\mathbb{E}[S_{\beta}^{1/\beta}]^{1-\delta}}}\right).$
@@
$E_{0}$ $=\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{\mathrm{d}}(\beta)}}\right)^{1/\beta}\mathopen{}\mathclose{{}\left(\log\frac{Z_{\mathrm{d}}(\beta)^{1/\beta}}{S_{\beta}^{1/\beta}}+\frac{1}{\beta^{\prime}}\log S_{\beta^{\prime}}-\frac{1}{\beta^{\prime}}\log Z_{\mathrm{d}}(\beta^{\prime})}\right)}\right]\eqqcolon E_{00}+E_{01}-E_{02},$
@@
$E_{02}$ $=\frac{1}{\beta^{\prime}}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{\log Z_{\mathrm{d}}(\beta^{\prime})}{Z_{\mathrm{d}}(\beta)^{1/\beta}}}\right]=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\log^{2}\mathopen{}\mathclose{{}\left(Z(\beta^{\prime})\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]}\right)}\right]^{1/2}\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-2/\beta}}\right]^{1/2}}\right),$
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]\geq\frac{1}{4}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\mathopen{}\mathclose{{}\left(1\wedge\frac{1}{R}}\right)}\right]=\frac{1}{4}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\wedge\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{\mathrm{d}}(\beta)}}\right)^{1/\beta}}\right].$ (3.12)
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]\geq\frac{1}{4M}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{\mathrm{d}}(\beta)}}\right)^{1/\beta}}\right]-\frac{1}{M^{\delta}}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{Z_{\mathrm{d}}(\beta^{\prime})}{S_{\beta^{\prime}}}}\right)^{\delta/\beta^{\prime}}\mathopen{}\mathclose{{}\left(\frac{S_{\beta}}{Z_{\mathrm{d}}(\beta)}}\right)^{(1+\delta)/\beta}}\right]}\right).$ (3.13)
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[Q_{\mathrm{d}}(\beta,\beta^{\prime})}\right]=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\frac{S_{\beta^{\prime}}}{Z_{\mathrm{d}}(\beta^{\prime})}}\right)^{1/\beta^{\prime}}\frac{1}{R^{1-\theta}}}\right]}\right)=O\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{(1-\theta)/\beta}S_{\beta^{\prime}}^{\theta/\beta^{\prime}}}\right]\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{\mathrm{d}}(\beta^{\prime})^{-\theta/\beta^{\prime}}}{Z_{\mathrm{d}}(\beta)^{(1-\theta)/\beta}}}\right]}\right),$ (3.14)
using the definition of $R$ and the independence of $(S_{\beta},S_{\beta^{\prime}})$ and $(Z_{\mathrm{d}}(\beta),Z_{\mathrm{d}}(\beta^{\prime}))$. By Cauchy–Schwarz inequality and Corollary 2.8,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{Z_{\mathrm{d}}(\beta^{\prime})^{-\theta/\beta^{\prime}}}{Z_{\mathrm{d}}(\beta)^{(1-\theta)/\beta}}}\right]\leq\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta^{\prime})^{-2\theta/\beta^{\prime}}}\right]^{1/2}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]^{\theta}}\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[Z(\beta)^{-2(1-\theta)/\beta}}\right]^{1/2}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{1-\theta}}=O\mathopen{}\mathclose{{}\left(\frac{(\beta-1)^{1-\theta}}{\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta^{\prime}}^{1/\beta^{\prime}}}\right]^{\theta}\mathbb{E}\mathopen{}\mathclose{{}\left[S_{\beta}^{1/\beta}}\right]^{1-\theta}}}\right),$